OpenStudy (anonymous):

Write the function f(x)= 2x^2+8x+7 in the form of f(x)= a(x-h)^2 + k and find the vertex?

OpenStudy (amistre64):

might want to complete the square, or expand it to compare parts

OpenStudy (anonymous):

im a bit confused on completing the square..

OpenStudy (amistre64):

then lets expand the right part of that to compare stuff with ...

OpenStudy (anonymous):

okay. so its 2(x^2+4)= -7?

OpenStudy (amistre64):

you forgot an x (and the equal sign is a bit odd), but yes, thats a decent approach to completeting the square 2x^2+8x+7 2(x^2+4x)+7 a complete square is a perfect square from some setup as: (x+n)^2 (x+n)^2 = x^2 +2nx + n^2 comparing this to what we have: x^2 + 4 x x^2 +2nx + n^2 notice that 2n = 4, allows us to find n^2

OpenStudy (amistre64):

if we +n^2 -n^2 to the original setup, then we are only adding 0 to it and nothing changes but the form.

OpenStudy (amistre64):

2n = 4, when n=2; and 2^2 = 4 sooo 2(x^2+4x +4 - 4)+7 2(x^2+4x +4) -2(4)+7 2(x+2)^2 -8+7 is this working out for us?

OpenStudy (anonymous):

yes so then the final form would be f(x)=2(x+2)^2-1 ?

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

so: f(x) = 2(x+2)^2 -1 do you know how to find the vertex from this?

OpenStudy (anonymous):

i believe so, i think it would be (-2, -1) ?

OpenStudy (amistre64):

thats correct

OpenStudy (anonymous):

thanks for your help!

OpenStudy (amistre64):

youre welcome